You are tasked to organize a speed dating event for your barangay. Based on previous events, the ratio of male to female attendees is 4:3. For your speed dating event to be successful, you need an equal ratio. 


If 24 males will attend the speed dating event, how many more females do you need to invite from other barangays to ensure an equal ratio of attendees?




1
Expert's answer
2021-01-24T17:28:39-0500

"\displaystyle
\textsf{Ratio} \,\, 4:3\\


\textsf{Let the number of attendants be}\,\, x\\

\frac{4}{7} \times x = 24\\

x = 42\\


\frac{3}{7} \times 42 = 18\\

\textsf{Difference}:\,\, 24 - 18 = 6.\\

\therefore 8\,\,\textsf{more remales are to be invited from other barangays}\\\textsf{to ensure an equal ratio of attendees.}"